Plainwell Polymers is going to producing the main dashboard structure for a new car design. The dashboard will be produced in three colors with expected daily demand as follows:

Black: 490 units/day

Grey: 145 units/day

Tan: 185 units/day

The injection-molding machine can produce a total of 1,055 parts per day running 24 hours per day. If Plainwell is going to produce the parts using a cyclic schedule and it takes 2 hours to change from one color to the next, what is the minimum cycle time for the cyclic schedule in hours? (Please show formula(s) to solve)
Answer #1

According to the problem

Total hours of working =24hours

Total minutes of working =24×60=1440

2hours to change from one color to next

2×60=120

Total productive mi utes=total minutes of working _(non productive)

1440_120=1320 minutes is productive

Idle time=toal productive ÷total total requirements

Black:1320÷490=2.7

Grey:1320÷145=9.1

Tan:1320÷185=7.1

Cycle time =no.of productive hours÷total units produced

24÷1055=.02 hour